OpenStudy (anonymous):

find the slope of the tangent line: y=7x^1/2+x^3/2, x=25

OpenStudy (anonymous):

first find the derivative of the line

OpenStudy (anonymous):

(7x^5/2+24)/(2x^4)

OpenStudy (anonymous):

what the...?

OpenStudy (anonymous):

my bad trying to do two problems at once...

OpenStudy (anonymous):

but the derivative is 82/10?

OpenStudy (anonymous):

i doubt it

OpenStudy (anonymous):

i didnt think so either but i cant find it

OpenStudy (anonymous):

\[y=7x^{1/2}+x^{3/2}\] is this your problem?

OpenStudy (anonymous):

correct and x=25

OpenStudy (anonymous):

dont care about the x=25 the derivative of x^n is n*x^(n-1)

OpenStudy (anonymous):

\[y'=\frac{7}{2}x^{-1/2}+\frac{3}{2}x^{1/2}\] plug in for x and solve for y' and thats the slope at x=25

OpenStudy (anonymous):

5.44341649?

OpenStudy (anonymous):

\[y= 7x^\left( \frac{ 1 }{ 2 } \right)+ x^\left( \frac{ 3 }{ 2 } \right) , x=25\] you mean this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[x^{1/2} =\sqrt{x}\] \[x^{-1/2}=\frac{1}{\sqrt{x}}\]

OpenStudy (anonymous):

\[\frac{ dy }{ dx }= \frac{ 7 }{2 }x^\left( \frac{ -1 }{ 2 } \right) + \frac{ 3 }{ 2 }x^\left( \frac{ 1 }{ 2 } \right) ; x=25\]

OpenStudy (anonymous):

i need the slope of the tangent line

OpenStudy (anonymous):

dy/dx is the slope...

OpenStudy (anonymous):

ya the slope would be \[\frac{ dy }{ dx }\]

OpenStudy (anonymous):

I keep plugging x=25 into f'(x) and getting my original m=82/10.

OpenStudy (anonymous):

it might be \[\frac{ 82 }{ 10 }\]

OpenStudy (anonymous):

i tried that and it says incorrect

OpenStudy (anonymous):

then either the question is wrong or you put the wrong question here ;)

OpenStudy (anonymous):

ya idk why its not working